# theoTheorem

### v2

@@ V Conclusion
This study has examined the impact of laser phase noise on adiabatic state preparation in a one-dimensional Rydberg spin chain, focusing on its role in driving excitations and influencing thermalization dynamics. By incorporating phase noise sampled from realistic laser power spectra, the simulations provide a realistic assessment of how coherence errors in the transverse field affect quantum state preparation. The results highlight a competition between diabatic and noise-induced excitations, leading to an optimal ramp duration that minimizes total energy absorption and improves final fidelities of the prepared state. The role of noise resonance and noise scaling is also examined, showing that phase noise is most disruptive when its dominant frequency components are resonant with key system excitations. Adjusting the spectral profile of the noise or shifting the system’s operational frequency away from the most affected regions improves state preparation fidelity. Additionally, the results demonstrate that increasing the overall noise strength leads to a rapid decline in performance, underscoring the need for noise suppression techniques in scalable neutral atom platforms.
@@
Excitation due to phase noise is found to depend strongly on the integrability of the system, with non-integrable regimes seen in the simulated adiabatic ramp facilitating energy delocalization, while integrable regimes lead to energy clustering and large spectral gaps which limit the energy transfer to low probability localized excitations. The interplay of these two regimes has direct implications on how energy is dissipated across the system, as we investigate whether this resembles many-body thermalization. Analysis of long-time expectation values indicates that noise-driven excitations can, under certain conditions, lead to thermalization consistent with the eigenstate thermalization hypothesis (ETH). The observed convergence between diagonal and thermal ensembles suggests that in these systems, the excitations generate correlation functions that are close to what would be expected in a thermal ensemble. However, we should note that in important parts of the parameter space, these systems are approximately integrable and will generally lead to deviations from thermalization – even more so when the noise strength is high enough to induce excessive heating. In contrast, in strictly integrable regimes, the absence of delocalized excitation channels restricts energy transfer across the spectrum and for higher energy states in particular, also preventing the system from reproducing thermal predictions.
@@
Our results have broad implications for quantum simulation and computation with Rydberg atom arrays, especially as system sizes increase and noise mitigation becomes a growing challenge. Future research in this topic could explore a broader range of phase noise spectra [24], higher-dimensional Rydberg geometries, different Rydberg accessible Hamiltonians, as well as observing the effects of phase noise on different interacting Hamiltonians. Also of interest is a parallel investigation of phase noise in the Van der Waals Rydberg interaction regime in which interactions scale as $1/r^{6}$ as opposed to the $1/r^{3}$ scaling of dipole-dipole interactions, and has become common in experimental setups in recent years [55, 56, 57, 58].
@@ Appendix A Noise Generation
The method used for noise generation is based on the TK95 algorithm [60] for the generation of independent time signals from a frequency power spectrum by applying a complex Gaussian filter followed by the inverse Fourier transform. To begin with, a sample spectral density of frequency noise fluctuations $S_{\nu}$ is chosen (see Fig. 3).d The frequency spectral density can be transformed into a phase spectral density by applying $S_{\nu}=\nu^{2}S_{\phi}$. In case a voltage power spectral density is provided $S_{V}$, we refer the reader to [61] where the conversion to $S_{\phi}$ is discussed in detail. To create a real noise signal, a double-sided conjugate symmetric spectrum must be generated. Given that the provided spectrum includes only positive frequencies, it is necessary to sample the spectrum twice using a complex Gaussian factor that is conjugated for negative frequencies. A frequency spectrum $S_{n}$ sampled from the phase noise power spectrum $S_{\phi}$ is defined by
@@
where the relevant units are shown in square brackets and $\mathcal{N}(0)$ is a Gaussian random number used to sample $S_{\phi}$ at each frequency ±$\pm k$ up until the Nyquist frequency $\nu_{Nyq}$. Once a trajectory of a phase noise spectrum $S_{n}$ is generated, the corresponding phase noise signal $\phi(t)$ is derived by an inverse Fourier transform
@@
where $\phi_{0}$ and $\phi_{\tau}$ are the same noise signal starting from $t=0$ and a lag of $t=\tau$ respectively, and $\sigma$ gives the standard deviation. Note that the numerator here is just the covariance between $\phi_{0}$ and $\phi_{\tau}$ where $E$ signifies the expected value of the given time series. Fig. 9 shows the agreement in autocorrelation functions of the original power spectrum $S_{\phi}$ and an average of 100 independently generated noise realizations.
Appendix B Excitation dynamics
@@
B.1 Reflection symmetry
@@
Figure 11: Above are plots of the diagonal ensemble as a function of energy eigenstates after a time evolution with four time independent Hamiltonians and laser phase noise. The ground state of the Hamiltonian from Eq. (II) is evolved with noise for a total time of $tV_{dd}=400$, with constant parameters $V_{dd}=1/2\pi$MHz, $\delta/V_{dd}=1.1$, and (a) $\Omega/V_{dd}=1.0$, (b) $\Omega/V_{dd}=0.6$, (c) $\Omega/V_{dd}=0.4$, (d) $\Omega/V_{dd}=0.2$.The plots are reproduced for two different noise scale factors $\kappa$ which move the peak power of the sampled noise PSD to higher and lower frequencies in Fourier space (see Fig. 5). In orange we have the diagonal ensemble for $\kappa=0.9$ which was shown to cause most excitation during the adiabatic state preparation, while in blue we have $\kappa=0.5$ corresponding to a PSD that is further from resonance with the system causing less excitation. Error bars plot standard error across 100 time evolutions with independent laser phase noise signals, but are too small to be seen suggesting consistent excitation pattern.
@@
The first two terms in the operator flip sites at either end of the chain while the third term ensures the preservation of the spin alignment if both sites are in the same state. The Hamiltonian can be shown to commute with this reflection operator, leading to two independent symmetry sectors of eigenstates made up of symmetric and anti-symmetric superpositions of computational states. For open boundary one-dimensional spin chains with an odd number of sites, there will always be computational states that are invariant under reflection. For example, the crystalline states discussed in Fig. 1 that dominate the ground state for low values of $\Omega$. Such reflection invariant computational states are only represented in the symmetric energy eigenstates. Furthermore, it can be shown that diabatic excitation, governed by changes to $\Omega$, as well as phase noise excitation, governed by changes in $\phi$, both respect this symmetry. This means that if all energy is initially in one of these symmetry sectors as is the case in the symmetric ground state at the beginning of the adiabatic state preparation, all dynamics will be confined to that symmetry sector. This effectively inhibits excitation in approximately half of the Hilbert space, thus halving the density of states that the noise can dissipate into.
Fig. 10 shows the separation of eigenenergies corresponding to symmetric eigenstates as a function of Rabi frequency $\Omega/V_{dd}$ and with constant detuning $\delta=1.1V_{dd}$, which are relevant parameters for the third step in the simulated adiabatic state preparation. Tracing the bottom lone which represents the ground state we see the minimum ground state energy gap corresponding to the critical region of the preparation that occurs between $\Omega=0.2-0.3V_{dd}$. Furthermore, the total energy span of the Hilbert space increases from $10V_{dd}$ to $6V_{dd}$ as $\Omega/V_{dd}$ increases. Initially, clusters of energy eigenvalues form an underlying structure of clustered energy eigenstates that is consistent with what one might expect in an Ising Hamiltonian, but as $\Omega/V_{dd}$ we see this structure break down as the clusters merge. This merging corresponds to the breaking of integrability that occurs when a transverse field is applied to the Ising Hamiltonian, which has direct implications for phase noise excitation.
B.2 Laser phase noise excitations in time independent Hamiltonians
While it is simple to treat the adiabatic state preparation as a black box and analyze only the final state, getting a deeper understanding of the complicated excitation dynamics of laser phase noise a time dependent Hamiltonian with ever changing energy level structure is much harder. Fig. 11 provides a small glimpse by choosing four different Hamiltonians at various values of $\Omega$ that occur throughout the simulated adiabatic state preparation, performing a time independent evolution of the ground state with laser phase noise. The resulting diagonal ensembles for two separate noise signals generated with different power profiles demonstrate the diminishing effects of laser phase noise as $\Omega$ is reduced and the system becomes increasingly integrable. The evolution of the ground state in the initial Hamiltonian of the preparation of the state in Fig. 11(a) results in the formation of well-defined equidistant peaks as the energy of the phase noise excitation cascades from one peak to the next. This excitation profile follows the matrix element predictions of Fig. 6(a) which plot the transition spectra of energy eigenstates corresponding to the four largest peaks, for which excitation dynamics is dominated by single delocalized transitions that excite the entire energy spectrum. Furthermore, although the excitation patterns for the two noise signals used in Fig. 11(a) are identical, the noise signal corresponding to a PSD rescaled by $\kappa=0.9$ known to result in the worst performance in state preparation (see Fig. 5(b)) results in a final state that has ground state fidelity of $\langle\psi_{gr}|\psi\rangle_{\kappa=0.9}=0.58$, compared to $\langle\psi_{gr}|\psi\rangle_{\kappa=0.5}=0.77$ for $\kappa=0.5$. In other words, the amount of phase noise excitation is highly dependent on the level of power a given realization has in the frequency region of the dominant transitions of the Hamiltonian. The Hamiltonian in Fig. 11(b) has $\Omega/V_{dd}=0.6$ which occurs near the middle of the simulated state preparation, and despite a contraction of the energy spectrum we still observe peaked features as for the initial Hamiltonian. This suggests that although the exact energy eigenstates might differ, the excitation dynamics in the initial stages of the state preparation exhibit similar behavior. However, upon closer inspection we see that the ground state fidelities have risen to $\langle\psi_{gr}|\psi\rangle_{\kappa=0.9}=0.67$ and $\langle\psi_{gr}|\psi\rangle_{\kappa=0.5}=0.81$, suggesting that the approach towards integrability is already affecting the total levels of excitation. Fig. 11(c) shows evolution with a Hamiltonian with $\Omega/V_{dd}=0.6$. We see a clear degradation of the peaked excitation spectra seen in the previous two plots, as the excitation dynamics shift towards more localized transition that lead to less overall energy mobility as the highest energy states are left largely unexcited. A drop in overall excitation is also evidenced by the rise in final ground state fidelities of $\langle\psi_{gr}|\psi\rangle_{\kappa=0.9}=0.84$ and $\langle\psi_{gr}|\psi\rangle_{\kappa=0.5}=0.82$, as we see that the low-frequency noise signal generated using $\kappa=0.5$ is now marginally more likely to excite the system. The transition to localized excitation is completed by the final stages of the Hamiltonian as seen in Figure Fig. 11(d) where evolution with a transverse field of $\Omega/V_{dd}=0.6$ leads to much lower levels of excitation as final ground state fidelities rise further to $\langle\psi_{gr}|\psi\rangle_{\kappa=0.9}=0.96$ and $\langle\psi_{gr}|\psi\rangle_{\kappa=0.5}=0.95$. Apart from the excitation out of the ground state itself, all phase noise excitation is localized to a small energy region surrounding the state with an exponential drop off in excitation as you get further away.

### v3

@@ V Conclusion
This study has examined the impact of laser phase noise on adiabatic state preparation in a one-dimensional Rydberg spin chain, focusing on its role in driving excitations and influencing thermalization dynamics. By incorporating phase noise sampled from realistic laser power spectra, the simulations provide a realistic assessment of how coherence errors in the transverse field affect quantum state preparation. The results highlight a competition between diabatic and noise-induced excitations, leading to an optimal ramp duration that minimizes total energy absorption and improves final fidelities. The role of noise resonance and noise scaling is also examined, showing that phase noise is most disruptive when its dominant frequency components are resonant with key system excitations. The results also demonstrate that increasing the overall noise strength leads to a rapid decline in performance, underscoring the need for noise suppression techniques in scalable neutral atom platforms. Additionally, shifting the system’s operational frequency away from the most affected regions by changing the interaction strength between sites improves the fidelity of the state preparation.
@@
Excitation due to phase noise is found to depend strongly on the integrability of the system, with non-integrable regimes seen in the simulated adiabatic ramp facilitating energy delocalization, while integrable regimes lead to energy clustering and large spectral gaps which limit the energy transfer to low probability localized excitations. The interplay of these two regimes has direct implications on how energy is dissipated across the system, as we investigate whether this resembles many-body thermalization. Analysis of long-time expectation values indicates that noise-driven excitations can, under certain conditions, lead to thermalization consistent with the eigenstate thermalization hypothesis (ETH). The observed convergence between diagonal and thermal ensembles suggests that in these systems the excitations generate correlation functions that are close to what would be expected in a thermal ensemble. However, we should note that in important parts of the parameter space, these systems are approximately integrable and will generally lead to deviations from thermalization – even more so when the noise strength is high enough to induce excessive heating. In contrast, in strictly integrable regimes, the absence of delocalized excitation channels restricts energy transfer across the spectrum, and for higher energy states in particular, also preventing the system from reproducing thermal predictions.
@@
Our results have broad implications for quantum simulation and computation with Rydberg atom arrays, especially as system sizes increase and noise mitigation becomes a growing challenge. Future research in this topic could explore a broader range of phase noise spectra [27], higher-dimensional Rydberg geometries, as well as observing the effects of phase noise on different Rydberg accessible interacting Hamiltonians. Also of interest is a parallel investigation of phase noise in the Van der Waals Rydberg interaction regime in which interactions scale as $1/r^{6}$ as opposed to the $1/r^{3}$ scaling of dipole-dipole interactions, and has become common in experimental setups in recent years [38, 64, 63, 14].
@@ Appendix A Noise Generation
The method used for noise generation is based on the TK95 algorithm [61] for the generation of independent time signals from a frequency power spectrum by applying a complex Gaussian filter followed by the inverse Fourier transform. To begin with, a sample spectral density of frequency noise fluctuations $S_{\nu}$ is chosen (see Figure 3).d The frequency spectral density can be transformed into a phase spectral density by applying $S_{\nu}=\nu^{2}S_{\phi}$. In case a voltage power spectral density is provided $S_{V}$, we refer the reader to [56] where the conversion to $S_{\phi}$ is discussed in detail. To create a real noise signal, a double-sided conjugate symmetric spectrum must be generated. Given that the provided spectrum includes only positive frequencies, it is necessary to sample the spectrum twice using a complex Gaussian factor that is conjugated for negative frequencies. A frequency spectrum $S_{n}$ sampled from the phase noise power spectrum $S_{\phi}$ is defined by
@@
where the relevant units are shown in square brackets and $\mathcal{N}(0)$ is a Gaussian random number used to sample $S_{\phi}$ at each frequency $\pm k$ up until the Nyquist frequency $\nu_{Nyq}$. Once a trajectory of a phase noise spectrum $S_{n}$ is generated, the corresponding phase noise signal $\phi(t)$ is derived by an inverse Fourier transform
@@
where $\phi_{0}$ and $\phi_{\tau}$ are the same noise signal starting from $t=0$ and a lag of $t=\tau$ respectively, and $\sigma$ gives the standard deviation. Note that the numerator here is just the covariance between $\phi_{0}$ and $\phi_{\tau}$ where $E$ signifies the expected value of the given time series. Figure 10 shows the agreement in autocorrelation functions of the original power spectrum $S_{\phi}$ and an average of 100 independently generated noise realizations.
Appendix B Rotating frame transformations in a Rydberg Hamiltonian with laser phase noise
The standard representation of laser phase noise in a Rydberg Hamiltonian involves a stochastic phase term coupled to the transverse drive. It is often desirable to transform the system into a frame where the Hamiltonian remains purely real. We begin with the Hamiltonian for a Rydberg atom array in the fixed-frequency rotating frame
$\hat{H}_{R}(t)=\frac{\Omega(t)}{2}\sum^{N}_{k}\left[e^{-i\phi(t)}\hat{\sigma}_{k}^{+}+e^{i\phi(t)}\hat{\sigma}_{k}^{-}\right]-\delta(t)\sum^{N}_{k}\hat{n}_{k}+\sum_{k<l}V_{kl}\hat{n}_{k}\hat{n}_{l},$ (14)
where $\hat{\sigma}_{k}^{+}=|1\rangle_{k}\langle 0|_{k}$ and $\hat{\sigma}_{k}^{-}=|0\rangle_{k}\langle 1|_{k}$ are the raising and lowering operators, $\hat{n}_{k}$ is the number operator counting Rydberg occupation, and $\phi(t)$ is a stochastic variable representing phase fluctuations. To eliminate the phase factor from the Rabi term, we define a time-dependent unitary transformation
$\hat{U}(t)=\exp\left(-i\phi(t)\sum^{N}_{k}\hat{n}_{k}\right),$ (15)
which effectively rotates the basis states to a frame that is in line with the phase jitter of the laser. Under a time-dependent unitary transformation, the effective Hamiltonian in the new frame, $\hat{H}^{\prime}$, is given by the relation
$\hat{H}^{\prime}=\hat{U}\hat{H}_{R}\hat{U}^{\dagger}+i\left(\frac{\partial\hat{U}}{\partial t}\right)\hat{U}^{\dagger}.$ (16)
Applying the first term in this transformation, we observe that the detuning $\delta(t)$ and the dipole-dipole interaction $V_{kl}$ commute with $\hat{U}$ because they are built using only $\hat{n}_{k}$. Consequently, these terms remain invariant in the new frame. However, the transverse coupling term is modified by the transformation of the raising and lowering operators, where $\hat{U}\hat{\sigma}_{k}^{+}\hat{U}^{\dagger}=\hat{\sigma}_{k}^{+}e^{i\phi(t)}$ and $\hat{U}\hat{\sigma}_{k}^{-}\hat{U}^{\dagger}=\hat{\sigma}_{k}^{-}e^{-i\phi(t)}$. Substituting these into the Rabi term of Eq. (14) yields
$\frac{\Omega(t)}{2}\sum_{k}\left(e^{-i\phi(t)}e^{i\phi(t)}\hat{\sigma}_{k}^{+}+e^{i\phi(t)}e^{-i\phi(t)}\hat{\sigma}_{k}^{-}\right)=\frac{\Omega(t)}{2}\sum^{N}_{k}\hat{\sigma}_{x}^{k}.$ (17)
This step effectively “unwinds” the phase noise from the transverse drive. The second part of the transformation law in Eq. (16) accounts for the “inertial” or “gauge” term arising from the explicit time-dependence of the unitary operator. Differentiating $\hat{U}$ with respect to time gives
$\frac{\partial\hat{U}}{\partial t}=-i\dot{\phi}(t)\left(\sum^{N}_{k}\hat{n}_{k}\right)\exp\left(-i\phi(t)\sum^{N}_{k}\hat{n}_{k}\right).$ (18)
Multiplying by $i$ and $\hat{U}^{\dagger}$ results in the addition of a longitudinal term $\dot{\phi}(t)\sum\hat{n}_{k}$ to the Hamiltonian. Physically, $\dot{\phi}(t)$ represents the instantaneous frequency fluctuation of the laser, and can thus be understood as a stochastic jitter in phase. Combining all components, the final transformed Hamiltonian is expressed as
$\hat{H}^{\prime}=\frac{\Omega(t)}{2}\sum^{N}_{k}\hat{\sigma}_{x}^{k}-\left[\delta(t)-\dot{\phi}(t)\right]\sum^{N}_{k}\hat{n}_{k}+\sum_{k<l}V_{kl}\hat{n}_{k}\hat{n}_{l}.$ (19)
This result provides a rigorous mathematical justification for treating laser phase noise as a stochastic fluctuation of the detuning. In the context of numerical modeling using the TK95 algorithm, this transformation allows the use of purely real-valued matrices.
Appendix C Excitation dynamics
@@
C.1 On integrability
One of the goals of this analysis will be to reveal how integrability and spectral structure influence the ability of phase noise to spread energy throughout the system. However, before beginning the discussion on excitation dynamics is useful to define what is meant by integrability in the context of the Rydberg chain Ising Hamiltonian we are simulating in this paper, as this term may often be used to cover a wide range of phenomena across different systems. In this work, integrability refers to the regime where the system’s dynamics are governed by a complete set of conserved quantities, resulting in a highly structured and ‘predictable’ energy spectrum. In the limit of $\Omega\to 0$, the Hamiltonian reduces to a Ising model where the number of Rydberg excitations is conserved, as without the transverse field there is no mechanism to flip the states of individual sites. Consequently, eigenstates are simply computational basis states, and the ground state is determined by the competition between $\delta$ and $V_{dd}$. In this integrable regime, the system is ‘stuck’ in a specific energy sector with different degenerate configurations with the same number of Rydberg excitations. As the transverse field $\Omega$ is increased, these conservation laws are broken, and the system transitions into a non-integrable, or chaotic, regime. Although the transition to a non-integrable regime is gradual and impossible to pinpoint precisely, it is marked by a break-down of degeneracy as $\Omega$ introduces state mixing and energy state repulsion as the degenerate computational states form superposition states with varying energies. In larger $\Omega$ the once isolated energy sectors merge into one sector with a high density of eigenstates that have distinct energies, making dynamics in such a system much more complex, as different couplings between states across the entire energy spectrum become accessible. Dynamics in such non-integrable systems becomes much more difficult to calculate especially at larger system sizes. This breakdown of integrability acts as a bridge, allowing phase noise to couple the ground state to a dense manifold of excited states that would otherwise be inaccessible in the highly structured and disconnected energy spectrum of the strictly integrable Hamiltonian.
@@ C.1 On integrability
Figure 11 shows the separation of eigenenergies corresponding to eigenstates as a function of Rabi frequency $\Omega/V_{dd}$ and with constant detuning $\delta=1.1V_{dd}$, which are relevant parameters for the third step in the simulated adiabatic state preparation. Tracing the bottom lone which represents the ground state we see the minimum ground state energy gap corresponding to the critical region of the preparation that occurs between $\Omega=0.2-0.3V_{dd}$. Furthermore, the total energy span of the Hilbert space increases from $10V_{dd}$ to $6V_{dd}$ as $\Omega/V_{dd}$ increases. Initially, clusters of energy eigenvalues form an underlying structure of clustered energy eigenstates that is consistent with what one might expect in an Ising Hamiltonian, but as $\Omega/V_{dd}$ we see this structure break down as the clusters merge. This merging corresponds to the breaking of integrability that occurs when a transverse field is applied to the Ising Hamiltonian, which has direct implications for phase noise excitation.
C.2 Reflection symmetry
@@
The first two terms in the operator flip sites at either end of the chain while the third term ensures the preservation of the spin alignment if both sites are in the same state. The Hamiltonian can be shown to commute with this reflection operator, leading to two independent symmetry sectors of eigenstates made up of symmetric and anti-symmetric superpositions of computational states. For open boundary one-dimensional spin chains with an odd number of sites, there will always be computational states that are invariant under reflection. For example, the crystalline states discussed in Figure 1 that dominate the ground state for low values of $\Omega$. Such reflection invariant computational states are only represented in the symmetric energy eigenstates. Furthermore, it can be shown that diabatic excitation, governed by changes to $\Omega$, as well as phase noise excitation, governed by changes in $\phi$, both respect this symmetry. This means that if all energy is initially in one of these symmetry sectors as is the case in the symmetric ground state at the beginning of the adiabatic state preparation, all dynamics will be confined to that symmetry sector. This effectively inhibits excitation in approximately half of the Hilbert space, thus halving the density of states that the noise can dissipate into. Consequently, Figure 11 only shows symmetric eigenstates relevant to our simulations.
C.3 Laser phase noise excitations in time independent Hamiltonians
Figure 12: The ground state of the Hamiltonian from Eq. (II) is evolved with noise for a total time of $TV_{dd}=300$, for different laser parameters (a) $\Omega/V_{dd}=0.6$ and $\delta/V_{dd}=-3.0$, (b) $\Omega/V_{dd}=0.6$ and $\delta/V_{dd}=-1.0$, (c) $\Omega/V_{dd}=0.6$ and $\delta/V_{dd}=1.1$, (d) $\Omega/V_{dd}=0.1$ and $\delta/V_{dd}=1.1$. In orange we have the diagonal ensemble after noisy evolution with site interaction strength set to $V_{dd}/2\pi=10$MHz, while in blue we have the same simulation reproduced for an interaction strength of $V_{dd}/2\pi=0.16$MHz. Error bars plot standard error across 100 time evolutions with independent laser phase noise signals using the same noise profile from Figure 3, but are too small to be seen on the log scale suggesting a consistent excitation pattern.
While it is simple to treat the adiabatic state preparation as a black box and analyze only the final state, getting a deeper understanding of the complicated excitation dynamics of laser phase noise a time dependent Hamiltonian with ever changing energy level structure is much harder. The transition rates discussed in Figure 6 provide a road map for a given source of energy to excite the system to higher many-body states. In order to bring these transition rate snapshots to life, in this section we perform an analysis of excitation dynamics using the same time-independent Hamiltonians.
[…]
Figure 12(a) shows the noisy evolution of the ground state of a Hamiltonian that occurs at the beginning of stage 2 of the adiabatic protocol. The resulting excitation is very low in the $V_{dd}/2\pi=10$MHz regime with a final ground state occupation of 98.2% at a standard error of 0.2%, but much larger in the $V_{dd}/2\pi=0.16$MHz regime in which ground state population dropped to 21.1% with standard error of 33%, suggesting that proximity of the dynamical frequency of the system to the 0.480.48MHz peak is a key factor for driving excitation even in the far detuned regime with a low density of states with high transition rates. Figure 12(b) with laser parameters from the middle of stage 2 paints a similar picture but with a modest increase in excitation across the spectrum as the energy gap between states decreases and more states become viable for excitation. The $V_{dd}/2\pi=10$MHz regime here has a final ground state population 89.489.4% at a standard error of 2%, while the ground state of the $V_{dd}/2\pi=0.16$MHz regime drops further to 12.112.1% with a standard error increasing to 44%. In particular, energy in the $V_{dd}/2\pi=0.16$MHz regime is much more evenly distributed throughout the energy spectrum than in the case of the $V_{dd}/2\pi=10$MHz regime, which has a much higher drop-off in occupation of higher energy states as most energy is still concentrated in the ground state and states with the highest ground state transition rate. As predicted in the matrix element analysis, Figure 12(c) that uses a parameters from the transition point between stages 2 and 3 results in the most overall excitation, with a ground state occupancy of 81.581.5% at a 33% standard error for the $V_{dd}/2\pi=10$MHz regime, and an almost fully depleted ground state occupancy of 5.15.1% at a 11% standard error for the $V_{dd}/2\pi=0.16$MHz regime. The evolution of the $V_{dd}/2\pi=10$MHz regime results in the formation of well-defined equidistant peaks as the energy of the phase noise occupancy cascades from one high transition rate to the next with a gradually decreasing probability. While these peaks are still visible in the $V_{dd}/2\pi=0.16$MHz regime, the comparatively stronger effect of phase noise leads to higher energy transfer with occupancy largely equilibrating across the entire spectrum. All three of these figures with a high transverse field at $\Omega/V_{dd}=0.6$ have excitation profiles dominated by transitions that are delocalized in energy, allowing noise to access the entire energy spectrum over time. Moreover, the rate of energy transfer improves as $\delta$ becomes increasingly positive and the energy required for transitions between many-body states with different numbers of atoms in a Rydberg state drops (see Figure 11). However, for the Hamiltonian that occurs at the end of the protocol seen in Figure 12(d), in which $\delta$ is positive but the transverse field drops down to $\Omega/V_{dd}=0.1$, there is a dramatic decrease in excitation across both regimes as the final ground state occupancy is 99.499.4% and 90.390.3% for the $V_{dd}/2\pi=10$MHz and $V_{dd}/2\pi=0.16$MHz regimes, respectively. The main difference is that the system here is much more integrable, leading to available transition rates that are not only lower but also highly localized in energy, meaning that phase noise transfers energy from the ground state at a much lower rate and is restricted to low energy sectors.